What is the wavelength (in nm) of the line in the spectrum of the hydrogen atom that arises from the transition of the electron from the orbital with n =6 to the orbital with n = 3?
Answer #1

Here photon will be emitted

1/lambda = R* (1/nf^2 - 1/ni^2)

R is Rydberg constant. R = 1.097*10^7

1/lambda = R* (1/nf^2 - 1/ni^2)

1/lambda = 1.097*10^7* (1/3^2 - 1/6^2)

lambda = 1.094*10^-6 m

lambda = 1094 nm

Answer: 1094 nm